if $ 1000 is invested now, $ 1500 two years from now, $ 2000 four years from now on interest rate of 6% compound annually, how much money accumulates after 10 years?
Answer by Theo(13342)

you are investing:

1000 for 10 years.
1500 for 8 years.
2000 for 6 years.

interest rate is 6% per year compounded annually.

formula is f = p * (1+r)^n

f is the future value.
p is the present value
r is the interest rate per time period (years in this case).
n is the number of time periods (years in this case).

first formula is f1 = 1000 * 1.06^10
second formula is f2 = 1500 * 1.06^8
third formula is f3 = 2000 * 1.06^6

solution is f1 + f2 + f3 which is equal to 1790.847697 + 2390.772112 + 2837.038225 = 7018.658033.